A 4 meter by 6 meter rug covers half the floor area of a room and leaves a uniform strip of base floor around the edges. That are the dimensions of the room?
(4*6)(2)=n I am not sure what the correct equation is for this problem.

Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(22740)

A 4 meter by 6 meter rug covers half the floor area of a room and leaves
a uniform strip of base floor around the edges.
That are the dimensions of the room?
:
Find the area of the rug: 4 * 6 = 24 sq/m
Find the area of the floor: 2 * 24 = 48 sq/m (twice the rug area)
:
Let x = the width of the uniform strip around the rug
:
Room length * room width = 48 sq./m
(2x+6) * (2x+4) = 48
FOIL
4x^2 + 8x + 12x + 24 = 48
:
4x^2 + 20x + 24 - 48 = 0
:
4x^2 + 20x - 24 = 0; our old friend the quadratic equatio16
:
Simplify, divide by 4:
x^2 + 5x - 6 = 0
:
Factor
(x + 6)(x - 1) = 0
:
x = +1 m is the width of the uniform strip, increases the dimensions by 2 m
therefore
6+2 = 8 m is the room length
and
4+2 = 6 m is the room width

Answer by Alan3354(69443)

A 4 meter by 6 meter rug covers half the floor area of a room and leaves a uniform strip of base floor around the edges. That are the dimensions of the room?
---------------
The run covers 24 sq meters.
The area (4+x)*(6+x) covers 2 times that, = 48 sq meters
x^2 + 10x + 24 = 48
x^2 + 10x - 24 = 0
(x+12)*(x-2) = 0
x = 2 meters
The room is 6 by 8 meters.